OpenStudy (anonymous):

y= sqrt 3x Domain Range y= sqrt (5x-3) Domain Range

OpenStudy (anonymous):

do you know the domain of \[y=\sqrt{x}\]

OpenStudy (anonymous):

basically, everything under the square root symbol has to be positive, so what does that mean for the domain?

OpenStudy (koikkara):

lol, isn't it... \(y=\sqrt{3x}\)

OpenStudy (yolomcswagginsggg):

y = sqrt 3x D : 3x ≥ 0 x ≥ 0 R : y ≥ 0 y = sqrt (5x - 3) D : 5x - 3 ≥ 0 x ≥ 3/5 R : y ≥ 0 lol ;p

OpenStudy (freckles):

it doesn't matter the sqrt(3) is a constant multiple the function's domain depends on sqrt(x)

OpenStudy (freckles):

\[f(x)=\sqrt{3x}=\sqrt{3} \sqrt{x} \\ \text{ the domain depends on the } \sqrt{x} \text{ for this function }\]